OpenStudy (akohl):

@Abhisar At the start of a race, car X, car Y, and car Z each accelerate from rest. Car X is more massive than car Y, and car Y is more massive than car Z. Over 10 seconds, the average force exerted on each car is identical. After 10 seconds, which car, if any, has the greatest momentum? Car X All have the same momentum Car Y Car Z

OpenStudy (abhisar):

What do you think ?

OpenStudy (akohl):

I fell like it should be the same since the force is the same

OpenStudy (abhisar):

Correct !

OpenStudy (abhisar):

Force = Rate of change of momentum = Final - Initial momentum/Time

OpenStudy (abhisar):

All the cars have same initial momentum = 0, let their final momenta be x, y and Z

OpenStudy (akohl):

Thanks that makes sense now

OpenStudy (abhisar):

F=(X-0)/10 , F=(Y-0)/10 and (F=z-0)/10 (wr F is the force which is similar in all cases.). Thus X=Y=Z

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)